# Linear-Size Ancilla Systems for Logical Measurements in QLDPC Codes

### v1

@@ Abstract
We show how to perform all logical Clifford gates on the [[144,12,12]] bivariate bicycle code, also known as the gross code. The scheme adds about 100 ancilla qubits into the connectivity graph, and one of the twelve logical qubits is sacrificed for gate synthesis. Logical measurements are combined with the automorphism gates studied by Bravyi et al. (Nature 627, 778-782) to implement 288 Pauli product measurements. We demonstrate the practicality of our scheme through circuit-level noise simulations, leveraging a novel modular decoder that combines BPOSD with matching.
The main technical contribution is a lower overhead logical measurement scheme based on gauge-fixing the construction by Cohen et al. (Sci. Adv. 8, eabn1717). Our techniques apply to general CSS codes and leverage expansion properties of the Tanner graph to give rigorous guarantees on qubit count, code distance, fault distance, and decoding distance of the modular decoder. In particular, we require $O(d/\beta)$ additional qubits where $\beta$ is the boundary Cheeger constant of the subgraph supporting the logical operator being measured. By introducing $O(d)$ additional bridge qubits, we are also able to measure products of logical Pauli operators and logical $Y$ operators.
@@ 1 Introduction
Reliable quantum computing at scale requires encoding quantum information into error-correcting codes and acting on that information without spreading errors too quickly [ABO08, Kit97, KLZ98, AGP05, Got09]. This necessitates a fault-tolerant architecture that both stores quantum information and also computes with it [CTV17]. The surface code provides the foundation for one of the leading fault-tolerant quantum computing architectures [BK98, DKLP02, FMMC12, Lit19]. It has a high accuracy threshold, allowing it to tolerate realistic noise rates, and its low-weight, geometrically local check operators can be measured on hardware that implements a two-dimensional lattice [KLR+22, AI23, SYK+23].
@@
One way forward is to relax the requirement of geometric locality, while retaining the need for qubits and checks to connect to only a few neighbors. Such high-rate low-density parity check (LDPC) codes can have much better parameters, both in terms of their encoding rate and their ability to correct errors [TZ14, PK21, BE21b, BE21a, PK22, LZ22]. Using these codes, fault-tolerant quantum computing becomes possible in principle with only constant space overhead in the asymptotic limit [Got14, FGL20].
@@
Recently, protocols for quantum memories based on LDPC codes have been discovered [BVC+17, NFB17, HB23, BCG+24, XAP+24]. Efficient memories could be used in concert with the surface code to reduce space overhead. In some instances, these memory protocols simultaneously have high thresholds and low overheads at practical code block sizes, making them attractive candidates for near-term demonstrations [BCG+24]. A growing body of research aims to find improved codes [SHR24, VXHB24], decoders [GCR24, WB24], and implementations for these quantum memories [VYL+23, PGP+24, BDS+24].
@@
Beyond storing information in high-rate quantum LDPC codes, computing directly on that information is an active area of research. There is the need for fast, parallel, and space-efficient gates, and also the challenge that logical qubits are encoded together in the same block, making them difficult to address individually. For stabilizer codes, there is a general solution that teleports qubits or gates and measures logical observables using specially prepared ancilla states [Got97, GC99, SI05, Kni05]. However, these large Steane- or Knill-type ancilla states are expensive to prepare, so new methods have been developed that are specific to LDPC codes [BVC+17, LB18, JO19, KP21, CKBB22, HJOY23, ZSP+23]. A subset of these approaches generalize the idea of lattice surgery to codes other than the surface code [BVC+17, LB18, KP21, CKBB22]. The methods can be complemented by symmetry-based gates, such as transversal and fold-transversal gates [QWV23, BB24, ES24].
@@
In this paper, we consider the problem of computing with LDPC codes using generalized lattice surgery. Our high level goal is to efficiently measure a set of logical Pauli operators of one or more LDPC code blocks. There are two simple, well-known ways to measure a weight $d$ logical operator: individually measure each qubit in the support of the logical operator and infer the logical qubit state, or entangle a single ancilla qubit to the operator using $d$ CNOT gates and measure the ancilla. It is easy to see that the first method can be highly disruptive to the remaining logical information, while the second method requires qubit degree $d+1$ and is not fault-tolerant, since an error on the ancilla can propagate to many other qubits. The problem of designing logical measurement schemes, therefore, is to obtain the measurement result fault-tolerantly while maintaining sparse connectivity and reasonable space overhead. We build on the merged codes and ancilla systems introduced by Cohen, Kim, Bartlett, and Brown [CKBB22], which are an application of stabilizer weight reduction techniques [Has21, SGI+24].
@@
We measure the quality of a logical measurement scheme in terms of four metrics: space-time overhead, parallelism, connectivity, and fault-tolerance. Let us now describe these metrics.
@@
    Space-time overhead: A scheme requires a system of ancilla qubits to implement a set of measurements. The total number of qubits in this system is an important metric. A circuit acting on the original code and the ancilla system enacts the measurement. The depth of this measurement circuit is another key metric.
@@
    Parallelism: Commuting logical operators can be measured simultaneously. A scheme allows some set of measurements to occur in parallel.
@@
The major goals of this paper are to significantly reduce the space overhead of the CKBB scheme and apply our results to a practically-motivated case study. Our scheme, which we name the gauged ancilla system, improves upon the CKBB scheme by gauge-fixing low-weight operators that may otherwise harm the distance of the merged code. We further improve the scheme by introducing bridge qubits to preserve distance during joint measurements. As we will show, these techniques, together with careful choice of what logical operators to measure, result in substantial improvement in space overhead.
Our central result is the proposed gauged ancilla system, which we put forward as a practical approach to fault-tolerantly measuring logical information in LDPC codes. We show that the number of layers in the ancilla system can be greatly reduced in general if the Tanner graph induced by the logical operator is suitably expanding; see Table 1. We further present proposals for gauged Y and joint measurement ancilla systems, as well as a modular decoding technique to reduce the decoding complexity during the logical measurement protocol. We systematically analyze the fault-tolerance properties of our proposals.
Ref [CKBB22] Theorem 6 in this paper
Ancilla system size $O(d^{2})$ $O(d/\beta)$
Bound on number of layers $L$ $L\geq 2d-1$ $\lceil L/2\rceil\geq 1/\beta$
Table 1: Parameters of merged codes with code distance $d$. The parameter $\beta$ is the boundary Cheeger constant of the Tanner graph induced by the logical observable.
Bringing our central results together, we present an in-depth study of a logical measurement scheme for the $[[144,12,12]]$ gross code, a 6-limited bivariate bicycle code [BCG+24]. In particular, we demonstrate that a mono-layer, degree-8 ancilla system with 103 additional qubits suffices to measure eight distinct logical Pauli operators, which significantly improves on the state of the art as summarized in Table 2. In this connectivity graph exactly two qubits have degree 8, and the remaining ones have degree 7 or less. We numerically study the logical error rates of $\bar{X}$ and $\bar{Z}$ logical measurements under circuit-level noise and find they are within a factor of 10x and 5x larger than that of the base code, respectively. We find that seven repetitions of the merged code syndrome measurement optimizes the logical and measurement error rates at a physical error rate of 0.001. We see that modular decoding gives comparable logical error rates to BPOSD but significantly improves decoding time. Finally, we combine permutation automorphisms with logical measurements to implement a set of 288 native Pauli measurements on the 12 logical qubits. This yields a further set of 95 Pauli rotations $\{\exp(i\frac{\pi}{4}P)\}$ that generate the eleven-qubit Clifford group within the gross code block. Futhermore the twelve logical qubits can be individually measured.
@@
The paper is organized as follows. Section 2 establishes notation and reviews the CKBB measurement scheme. Section 3 presents our low-overhead gauged ancilla systems for CSS codes, describes our modular decoding technique, and establishes the fault-tolerance properties of the systems and measurement protocols. In Section 4, we apply these techniques to construct an ancilla system for the $[[144,12,12]]$ gross code, numerically investigate its performance in the circuit model, and synthesize logical Clifford gates on the code block.
@@ 2.1 Tanner graph notation
This way of drawing the Tanner graph is essentially a description of the code as a chain complex. We can therefore think of sets of checks $C$ and sets of qubits $V$ as vector spaces over $\mathbb{F}_{2}$ with $F$ the boundary map between them. A subset $c\subset C$ can be seen as a row vector in $\mathbb{F}_{2}^{|C|}$. Then, if $F:C\to_{Z}V$, the qubits in $V$ that correspond to the product of the subset of checks $c$ in $C$ are $v=cF\in\mathbb{F}_{2}^{|V|}$.
@@
Similarly, we indicate a product of checks by $\mathcal{H}(c\in C)=\mathcal{H}(c)$ and the product of all checks in $C$ by $\mathcal{H}(C)$. We also use $\mathcal{H}_{Z}(\cdot)$ or $\mathcal{H}_{X}(\cdot)$ if the checks are entirely $Z$- or $X$-type, respectively. This is meant to indicate the Pauli operator obtained by multiplying together the checks indicated by vector $c$. For example, if $F:C\to_{Z}V$, then $\mathcal{H}_{Z}(c\in C)=Z(cF\in V)$.
@@ 2.2 The CKBB ancilla system
The number of ancilla qubits needed is $(2d-1)(|V_{0}|+|C_{0}|)>2d^{2}$, which is a daunting space overhead. For codes where the distance $d$ scales as $\sqrt{n}$, this number of ancilla qubits may be more than the number of qubits used in the code itself. One of the primary goals of this paper is to reduce this space overhead by using fewer layers. The time overhead is $d$ rounds of error correction which is standard albeit on a bigger code. We observe that if $\mathcal{G}$ has maximum degree $D$, then the CKBB ancilla system has maximum degree $D+1$.
@@
The CKBB ancilla system is a subsystem code with all the logical qubits from the original code $\mathcal{G}$ except the one being measured, and some gauge qubits which carry no logical information. It was proven in [CKBB22] that the CKBB ancilla system has code distance $d$. The fault distance or circuit distance of such a measurement protocol is not known.
3 Gauged Ancilla Systems
In this section we present our main result: a family of lower-overhead ancilla systems that measure logical operators in CSS codes. In sections 3.1 and 3.2 we detail the gauged $X$ ancilla system $\mathcal{G}_{X}$, which is similar to the CKBB $X$ ancilla described above but introduces additional gauge checks. In section 3.3 we prove Theorem 6, which relates the distance of the merged quantum code to the ancilla system size and the Cheeger constant of its induced Tanner graph. We introduce a modular decoding architecture in section 3.4. In sections 3.5 and 3.6 we generalize the construction to multi-qubit measurements $\mathcal{G}_{XX}$ (including between separate codeblocks) and $Y$ measurements $\mathcal{G}_{Y}$, which require an additional bridge construction.
3.1 Gauged $X$ Ancilla Systems

### v3

# Improved QLDPC Surgery: Logical Measurements and Bridging Codes

@@ Abstract
In this paper, we introduce the gauge-fixed QLDPC surgery scheme, an improved logical measurement scheme based on the construction of Cohen et al. (Sci. Adv. 8, eabn1717). Our scheme leverages expansion properties of the Tanner graph to substantially reduce the space overhead of QLDPC surgery. In certain cases, we only require $\Theta(w)$ ancilla qubits to fault-tolerantly measure a weight $w$ logical operator. We provide rigorous analysis for the code distance and fault distance of our schemes, and present a modular decoding algorithm that achieves maximal fault-distance. We further introduce a bridge system to facilitate fault-tolerant joint measurements of logical operators. Augmented by this bridge construction, our scheme can connect different families of QLDPC codes into one universal computing architecture.
Applying our toolbox, we show how to perform all logical Clifford gates on the [[144,12,12]] bivariate bicycle code. Our scheme adds 103 ancilla qubits into the connectivity graph, and one of the twelve logical qubits is used as an ancilla for gate synthesis. Logical measurements are combined with the automorphism gates studied by Bravyi et al. (Nature 627, 778-782) to implement 288 Pauli product measurements. We demonstrate the practicality of our scheme through circuit-level noise simulations, leveraging our proposed modular decoder that combines BPOSD with matching.
@@ 1 Introduction
Reliable quantum computing at scale requires encoding quantum information into error-correcting codes and acting on that information without spreading errors too quickly [ABO97, Kit97, KLZ98, AGP06, Got09]. This necessitates a fault-tolerant architecture that both stores quantum information and also computes with it [CTV17]. The surface code provides the foundation for one of the leading fault-tolerant quantum computing architectures [BK98, DKLP02, FMMC12, Lit19]. It has a high accuracy threshold, allowing it to tolerate realistic noise rates, and its low-weight, geometrically local check operators can be measured on hardware that implements a two-dimensional lattice [KLR+22, AI23, SYK+23].
@@
One way forward is to relax the requirement of geometric locality, while retaining the need for qubits and checks to connect to only a few neighbors. Such high-rate low-density parity check (LDPC) codes can have much better parameters, both in terms of their encoding rate and their ability to correct errors [TZ14, PK21, BE21b, BE21a, PK22, LZ22, DHLV23]. Using these codes, fault-tolerant quantum computing becomes possible in principle with only constant space overhead in the asymptotic limit [Got14, FGL18].
@@
Recently, several protocols for quantum memories based on LDPC codes have been developed [BVC+17, PNFB17, HB23, BCG+24, XBAP+24]. Efficient memories could be used in concert with the surface code to reduce space overhead. In some instances, these memory protocols simultaneously have high thresholds and low overheads at practical code block sizes, making them attractive candidates for near-term demonstrations [BCG+24]. A growing body of research aims to find improved codes [SHR24, VXHB24], decoders [GCR24, WB24], and implementations for these quantum memories [VYL+23, PGP+24, BDS+24].
@@
Beyond storing information in high-rate quantum LDPC codes, computing directly on that information is an active area of research. There is the need for fast, parallel, and space-efficient gates, and also the challenge that logical qubits are encoded together in the same block, making them difficult to address individually. For stabilizer codes, there is a general solution that teleports qubits or gates and measures logical observables using specially prepared ancilla states [Got97, GC99, SI05, Kni05]. However, these large Steane- or Knill-type ancilla states are expensive to prepare, so new techniques for logical computation on LDPC codes have been developed [BVC+17, LB18, JO19, KP21, CKBB22, HJOY23, ZSP+23]. One of these approaches generalizes the idea of lattice surgery to LDPC codes [BVC+17, LB18, KP21, CKBB22].
@@
In this paper, we consider the problem of computing with LDPC codes using generalized lattice surgery. Our high level goal is to efficiently measure a set of logical Pauli operators of one or more LDPC code blocks. Using Pauli-based computation [BSS16, Lit19], logical Pauli measurements can be used to implement Clifford gates and, provided additional magic states [BK05], non-Clifford gates. There are two simple, well-known ways to measure a weight $d$ logical operator: individually measure each qubit in the support of the logical operator and infer the logical qubit state, or entangle a single ancilla qubit to the operator using $d$ CNOT gates and measure the ancilla. It is easy to see that the first method can be highly disruptive to the remaining logical information, while the second method requires qubit degree $d$ and is not fault-tolerant, since an error on the ancilla can propagate to many other qubits. The problem of designing logical measurement schemes, therefore, is to obtain the measurement result fault-tolerantly while maintaining sparse connectivity and reasonable space-time overhead.
@@
We measure the quality of a logical measurement scheme in terms of four metrics: space-time overhead, parallelism and applicability, connectivity, and fault-tolerance. Let us now detail these metrics.
@@
    Space-time overhead: A scheme makes use of a system of ancilla qubits to implement a set of measurements. The total number of qubits in this system is an important metric. A circuit acting on the original code and the ancilla system enacts the measurement, so the depth of this measurement circuit is another key metric.
@@
    Parallelism and applicability: A scheme should be applicable to various settings, including single-logical qubit measurements, multi-logical qubit joint measurements, and joint measurements across multiple code blocks. A scheme should also support parallel measurements of commuting logical operators.
@@
In an earlier work [CKBB22], Cohen, Kim, Bartlett, and Brown introduced a scheme based on an ancilla system which is a hypergraph product code. We henceforth refer to their scheme as the CKBB scheme and explain its mechanism in Section 2.2. For a logical operator of weight $d$, the CKBB ancilla system has $12d-$ 1 layers, each layer isomorphic to the Tanner graph of the measured operator, totaling to a space overhead of $O(d^{2})$. This ancilla system can be naturally extended to support joint measurements. The scheme then use $d$ rounds of QEC on the merged code to obtain the logical measurement result. If the original code has qubit degree $\delta$ and distance $d$, the merged code will have degree $\delta+1$ and distance $d$. While the CKBB scheme has many desirable properties, its space overhead is daunting: for typical LDPC codes where the distance scales as $d\sim\Omega(\sqrt{n})$, the ancilla system needed for measuring one logical operator may be larger than the code block itself. This poses a significant barrier to utilizing the CKBB scheme in practice.
In this paper, we propose gauge-fixed QLDPC surgery as an improved version of the CKBB scheme with significantly lower space overhead. Our main theoretical contributions can be summarized as follows.
  1. 1.
    We gauge-fix the CKBB ancilla system so that the merged code contains no gauge qubits. The resulting gauge-fixed ancilla system ensures that the merged code has distance at least $d$, while using much fewer than $12d-$ 1 layers if the Tanner graph induced by the logical operator is expanding (Theorem 6). In cases where the graph is sufficiently expanding, the space overhead for measuring a weight $w$ logical operator is just $\Theta(w)$.
  2. 2.
    As a result of reducing the number of layers, the constructions of joint measurement systems from [CKBB22] are no longer distance preserving. We propose new, distance preserving constructions of $Y$ and joint measurement systems, utilizing a set of bridge qubits (Section 3.6 and 3.7).
  3. 3.
    We show that our gauge-fixed system with bridge can be used as an code adapter, which can fault-tolerantly connect two logical operators from two arbitrary code blocks (Section 3.8). This paves a path towards large-scale architectures based on multiple families of QLDPC codes.
  4. 4.
    We present a rigorous analysis of fault distance, and prove that $d$ rounds of syndrome extraction ensures fault distance $d$ for all of our proposed measurement schemes (Theorem 11,17,22), including the code adapter (Theorem 24).
  5. 5.
    Utilizing such an ancilla system increases the decoding complexity of the merged code. We propose a modular decoding algorithm on the $d$-round space-time decoding graph of the merged code (Section 3.5) which decomposes the decoding graph into two parts, one supported on the original code, the other supported on the gauge-fixed ancilla system, and decodes them separately. We prove that this modular decoder achieves fault distance $d$.
Our gauge-fixed surgery scheme comes with a noteworthy caveat: without additional assumptions on the measured operator or the code, we do not have upper bounds on the weight of the gauge-fixed operators. In other words, gauge-fixing these operators may break the LDPC property on the merged code. We discuss this caveat at length in Section 3.4, where we show theoretical evidence (Lemma 12) suggesting that for many important families of QLDPC codes, these gauge checks may not break the LDPC property. Notably, we prove that for hypergraph product codes the merged code has degree $\delta+1$ if the original code has degree $\delta$ (Theorem 13). Our bridge construction also lacks a worst case guarantee on the LDPC property, which we discuss in Section 3.8. We remark that after this work first appeared, these caveats are fully resolved in the schemes proposed by [WY24] and [SJOY24]. See section 1.1 for a more detailed discussion on related works.
Our theoretical analysis offers strong support for our practical optimizations and numerical simulations on the $[[144,12,12]]$ gross code, a 6-limited bivariate bicycle code [BCG+24]. Our main practical contributions are as follows.
  1. 1.
    We demonstrate that a mono-layer ancilla system with 103 additional qubits suffices to fault-tolerantly measure eight distinct logical Pauli operators, which significantly improves on the state of the art as summarized in Table 1. We verified numerically with CPLEX that all eight measurement protocols have fault distance $12$.
  2. 2.
    In the connectivity graph of the merged code there are exactly two qubits with degree 8, and the remaining qubits have degree 7 or less. The use of two degree-8 qubits is not a side effect of gauge-checks, but instead a result of using one ancilla system for multiple distinct Pauli measurements. For details, please see Figure 7.
  3. 3.
    We implemented a modular decoder using a combination of BPOSD and matching, and find that it gives comparable logical error rates to a fully BPOSD decoder with significantly faster decoding time (Figure 11).
  4. 4.
    Using the modular decoder, we numerically study the logical error rates of $\bar{X}$ and $\bar{Z}$ logical measurements under circuit-level noise and find they are within a factor of 10x and 5x larger than that of the base code, respectively (Figure 10). The measurement error rate is approximately 10x lower than the logical error rate, and we find that 7 repetitions of the merged code syndrome measurement optimizes the logical and measurement error rates at a physical error rate of 0.001.
  5. 5.
    Finally, we combine permutation automorphisms of the gross code with logical measurements to implement a set of 288 native Pauli measurements on the 12 logical qubits. These measurements yield a further set of 95 Pauli rotations $\{\exp(i\frac{\pi}{4}P)\}$ that generate the eleven-qubit Clifford group on the gross code block. Furthermore the twelve logical qubits can be individually measured.
@@
This practical progress reshapes our vision of a medium-to-large scale computing architecture using bivariate bicycle codes. We further remark that most of our results on the gross code are better than our worst case theoretical analysis. In particular, the fact that a mono-layer system is distance-preserving for all eight measurement protocols is a pleasant surprise. For this reason, to apply the gauge-fixed surgery scheme to other families of QLDPC codes, we suggest using our theoretical analysis as an “instruction manual”, and expect case-by-case optimized results to be better.
The paper is organized as follows. Section 2 establishes notation and reviews the CKBB measurement scheme. Section 3 presents our gauge-fixed surgery scheme for CSS codes and all aforementioned theoretical analysis. In Section 4, we apply our theoretical toolbox to construct an ancilla system for the $[[144,12,12]]$ gross code, numerically investigate its performance in the circuit model, and synthesize logical Clifford gates on the code block.
1.1 Related work
Since this work first appeared, a few others have followed with refinements to the central ideas. First, while we point out here that expansion existing in the Tanner graph of the code can drastically reduce the space-overhead of QLDPC surgery, the techniques of Refs. [WY24] and [IGND24] can simply add the required expansion to the ancilla system to achieve the same or better space overhead. In addition, Refs. [WY24, IGND24] think of the ancilla system as built on a graph and this perspective has other benefits in that certain cellulation [Has21, WY24, IGND24] and decongestion [FH21, Has21, WY24] ideas can be applied to guarantee that the gauge-fixed ancilla system remains LDPC. Second, while we point out here that any two ancilla systems can be joined via an adapter to measure a joint logical operator, Ref. [SJOY24] constructs the adapter in such a way to guarantee the newly introduced checks are also still LDPC. This construction relies also on the graphical perspective from Refs. [WY24, IGND24]. Third, both Refs. [XZZ+24] and [ZL24] present approaches to parallelizing related logical operations and appeared around the same time as our work. Ref. [XZZ+24] does this specifically for homological product codes using the similar technique of homomorphic measurements [HJOY23], while Ref. [ZL24] creates some parallelized surgery methods in the CKBB [CKBB22] style without gauge-fixing.
The independent work of [Cow24] appeared shortly before ours, in which QLDPC surgery is studied in the language of chain complexes and homology. [Cow24] further applied the CKBB construction to various small-scale QLDPC codes, and presented numerical results showing that in concrete instances, the space-overhead is often much less than the $O(d^{2})$ upper bound, which is reflected here in Table 1.
@@ 2.1 Tanner graph notation
This way of drawing the Tanner graph is essentially a description of the code as a chain complex. We can therefore think of sets of checks $C$ and sets of qubits $V$ as vector spaces over $\mathbb{F}_{2}$ with $F$ the boundary map between them. A subset $c\subset C$ can be seen as a row vector in $\mathbb{F}_{2}^{|C|}$. Then, if $F:C\to_{Z}V$, the qubits in $V$ that correspond to the product of the subset of checks $c$ in $C$ are $v=cF\in\mathbb{F}_{2}^{|V|}$. Due to this vector space interpretation, for the rest of this paper we will often write $c\in C$ to denote the subset indicated by $c$, and similarly for $v\in V$.
@@
In this paper, we frequently discuss the merged code obtained by attaching an ancilla system to the original code being measured. To differentiate between the two codes, we use $H$ to denote the stabilizer check matrices in the original code and use $\mathcal{H}$ for the merged code. For a set of checks $c\in C$, we denote their product by $\mathcal{H}(c\in C)=\mathcal{H}(c)$, and the product of all checks in $C$ by $\mathcal{H}(C)$. We also use $\mathcal{H}^{Z}(\cdot)$ or $\mathcal{H}^{X}(\cdot)$ if the checks are entirely $Z$- or $X$-type, respectively. This is meant to indicate the Pauli operator obtained by multiplying together the checks indicated by vector $c$.
@@ 2.2 The CKBB ancilla system
The number of ancilla qubits needed is $(2d-1)(|V_{0}|+|C_{0}|)>2d^{2}$, which is a daunting space overhead. For codes where the distance $d$ scales as $\sqrt{n}$, this number of ancilla qubits may be more than the number of qubits used in the code itself. One of the primary goals of this paper is to reduce this space overhead by using fewer layers. The time overhead is $d$ rounds of error correction which is standard albeit on a bigger code. We observe that if $\mathcal{G}$ has maximum degree $\delta$, then the CKBB ancilla system has maximum degree $\delta+1$.
@@
The CKBB ancilla system is a subsystem code with all the logical qubits from the original code $\mathcal{G}$ except the one being measured, and some gauge qubits which carry no logical information. It was proven in [CKBB22] that the CKBB ancilla system has code distance $d$. The fault distance or circuit distance of such a measurement protocol is not analyzed.
3 Gauge-Fixed QLDPC Surgery
In this section we present our main theoretical result: the gauge-fixed QLDPC surgery scheme. In Section 3.1 we introduce the basic form of the gauge-fixed ancilla system, under the context of measuring a $X$ logical operator. In Section 3.2 we detail the $X$ logical measurement protocol and space-time detector graph. We prove the code distance of the merged code, as well as the fault distance of the space-time detector graph in Section 3.3. We analyze the check weights of the gauge-fixed ancilla system in Section 3.4, and introduce the modular decoder in Section 3.5. In Sections 3.6 and 3.7 we generalize the construction to multi-qubit measurements and $Y$ measurements, which require an additional bridge system. In Section 3.8, we discuss how to use our bridge construction as a code adapter and connect logical operators from different families of codes. We prove the code and fault distances of these measurements in their respective sections.
3.1 Gauge-fixed $X$ Ancilla Systems